### v1

# Robustness of Interferometric Power to Sudden Death

Abstract
We study the dissipative dynamics of interferometric power as a discordlike measure in Markovian environments, such as dephasing, depolarizing, and generalized amplitude damping. Moreover, we compare the dynamics of interferometric power and entanglement by choosing proper initial conditions. Our study shows that in all cases where the sudden death of entanglement appears, interferometric power decays asymptotically. Therefore, quantum metrology based on interferometric power is more robust than entanglement.
keywords:
sudden death; entanglement; interferometric power.
1 Introduction
Entanglement and quantum discord are two kinds of quantum correlations that differ from classical correlations.[1, 2, 3, 4, 5] While entanglement is recognized as the main reason for the quantum computation advantage over classical computation, quantum discord found in separable mixed states plays a key role in quantum computing and quantum information processing. For instance, quantum discord can be used in study of application in deterministic quantum computation with one pure qubit (DQC1),[6] quantum phase transmission,[7, 8] and application of quantum correlations in the Grover search algorithm.[9]
Quantum interferometric power (IP), proposed in the context of quantum metrology, was first introduced by Girolami et al.[10] They proved that this quantity could be a rigorous measure of discord-type quantum correlations of an arbitrary bipartite state. As a discordlike quantum correlation measure, IP has several basic properties, such as nonnegativity, local unitary invariance and contractivity.[11] Furthermore, IP exists the phenomenon of the sudden change for a two-qubit X shape states under several different kinds of quantum noises[12], just like quantum discord.[13, 14, 15, 16]
When considering the dynamics of a two-qubit entangled state under noisy environments, disentanglement can occur in a finite time.[17, 18, 19, 20, 21, 22, 23, 24, 25] This phenomenon, so-called ”entanglement sudden death” (ESD), is different from the usual decoherence in asymptotic time. The occurrence of such a phenomenon depends on the initial two-qubit state and the system-environment interactions. Here we investigate the dynamics of IP of a two-qubit system under several different kinds of noisy channels where the ESD can occur in the same case. Our results show that IP decays only in the asymptotic limit when entanglement exhibits sudden death under the same noisy environments. That is, IP is more robust than entanglement in resisting decoherence, which implies that quantum metrology based on IP is more robust than based on entanglement.
2 The quantification of entanglement and quantum discord
Here, to investigate the dynamics of two-qubit states we use concurrence[26] and interferometric power[10] as the quantification of entanglement and quantum discord, respectively.
The concurrence of a two-qubit state $\rho_{AB}$ can be given as $C(\rho_{AB})=\max\{0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4}\}$, where $\lambda_{i}$s are the eigenvalues of the Hermitian matrix $\sqrt{\rho_{AB}\sigma_{y}\otimes\sigma_{y}\rho_{AB}^{\ast}\sigma_{y}\otimes\sigma_{y}}$, $\rho_{AB}^{\ast}$ being the complex conjugate of $\rho_{AB}$. The structure of density matrices we use to calculate concurrence has an $X$-shape, which means that the form of concurrence can be rewritten as
$C(\rho_{AB})=2\max\{0,\Lambda_{1},\Lambda_{2}\},$ (1)
where $\Lambda_{1}=|\rho_{14}|-\sqrt{\rho_{22}\rho_{33}}$, $\Lambda_{2}=|\rho_{23}|-\sqrt{\rho_{11}\rho_{44}}$.
The definition of IP for $\rho_{AB}$ with subsystem A being a qubit is[14]
$\mathcal{IP}^{A}(\rho_{AB})=\zeta_{min}[M],$ (2)
where $\zeta_{min}[M]$ is the smallest eigenvalues of the matrix $M$ with elements
$M_{m,n}=\frac{1}{2}\sum_{i,l;q_{i}+q_{l}\neq 0}\frac{(q_{i}-q_{l})^{2}}{q_{i}+q_{l}}\langle\psi_{i}|\sigma_{m}\otimes\mathbb{1}|\psi_{l}\rangle\langle\psi_{l}|\sigma_{n}\otimes\mathbb{1}|\psi_{i}\rangle$ (3)
with $\{q_{i},|\psi_{i}\rangle\}$ being the eigenvalues and eigenvectors of $\rho_{AB}$. By this definition, $\mathcal{IP}^{A}(\rho_{AB})$ can be considered as an operational and computable indicator of quantum discord.
3 Dynamics of concurrence and IP under noisy environments
3.1 Phase damping
Phase damping describes the loss of information without loss of energy. To investigate the dynamics of two-qubit entanglement and IP under dephasing noisy channel, we consider the initial state as Werner state $\rho_{W}(0)=(1-\alpha)\mathbb{1}/4+\alpha|\Psi^{-}\rangle\langle\Psi^{-}|$, where $\alpha\in[0,1]$ and $|\Psi^{-}\rangle=(|01\rangle-|10\rangle)/\sqrt{2}$. The Kraus operators for a dephasing channel we considered can be given by[1]
$E_{0}=\left(\begin{array}[]{cc}1&0\\ 0&\sqrt{1-\gamma}\end{array}\right),\ \ E_{1}=\left(\begin{array}[]{cc}0&0\\ 0&\sqrt{\gamma}\end{array}\right),$ (4)
where $\gamma=1-e^{-\Gamma t}$ and $\Gamma$ denotes transversal decay rate. During the evolution, the elements of the density matrix for output states are
$\begin{split}\rho_{ii}(t)&=\rho_{ii}(0),\ \ \ i=1,\dots,4\\ \rho_{23}(t)&=\rho_{32}(t)=\rho_{23}(0)(1-\gamma).\end{split}$ (5)
The explicit expression of concurrence for this state is $C(\rho)=\alpha(3/2-\gamma)-1/2$. It is easy to find that for any $\alpha\neq 1$, the concurrence will reach to zero in a finite time, as shown in Fig. 1. On the other hand, the expression of IP for this state is $\mathcal{IP}(\rho)=\min\{\alpha^{2}(2+\gamma)^{2}/[2(1+\alpha-\alpha\gamma)]\!+\!\alpha^{2}\gamma^{2}/[2(1-\alpha+\alpha\gamma)],2\alpha^{2}(1-\gamma)^{2}/(1+\alpha)\}$. One can see that for any $\alpha\neq 1$, the IP of $\rho(t)$ vanishes only in the asymptotic limit, as shown in Fig. 1.
Figure 1: Dynamics of (a) concurrence and (b) IP as functions of $\alpha$ and $\gamma$, under phase damping channel.
3.2 Generalized amplitude damping
Generalized amplitude damping (GAD) describes the effect of dissipation to an environment at finite temperature. The Kraus operators of this kind of noisy channel are[1]
$\begin{split}E_{0}&=\sqrt{q}\left(\begin{array}[]{cc}1&0\\ 0&\sqrt{1-\gamma}\end{array}\right),\ \ E_{1}=\sqrt{q}\left(\begin{array}[]{cc}0&\sqrt{\gamma}\\ 0&0\end{array}\right),\\ E_{2}&=\sqrt{1-q}\left(\begin{array}[]{cc}\sqrt{1-\gamma}&0\\ 0&1\end{array}\right),\ \ E_{3}=\sqrt{1-q}\left(\begin{array}[]{cc}0&0\\ \sqrt{\gamma}&0\end{array}\right),\end{split}$ (6)
where $q$ defines the probability distribution of the qubit when $t\rightarrow\infty$ and $\gamma$ is defined above.
Figure 2: Dynamics of (a) and (c) concurrence and (b) and (d) IP as functions of $\alpha$ and $\gamma$, under generalized amplitude damping channel.
Here we consider the initial state as $\rho(0)=|\Phi\rangle\langle\Phi|$ with
$|\Phi\rangle=\sqrt{1-\alpha}|00\rangle+\sqrt{\alpha}|11\rangle,\ \ \ \ \alpha\in[0,1].$ (7)
According to Eq. , the output density matrix under the generalized amplitude damping channel can be obtained,
$\begin{split}\rho_{11}(t)=(1-\alpha)\{1&-\gamma[2(1-q)-\gamma(1-2q)]\}+\gamma^{2}q^{2},\\ \rho_{22}(t)=\rho_{33}(t)=\gamma[&(1-\gamma)(1-2q)(1-\gamma)+q(1-\gamma q)],\\ \rho_{44}(t)&=1-\rho_{11}(t)-2\rho_{22}(t),\\ \rho_{14}(t)=&\rho_{41}(t)=(1-\gamma)\sqrt{\alpha(1-\alpha)}.\end{split}$ (8)
Using formula (1-3), the concurrence and the IP of this output state can be obtained as $C_{GAD}[\rho(t)]=2\max\{0,\Lambda_{1}(t)\}$ and $\mathcal{IP}^{A}[\rho(t)]\!=\!\min\{M_{11},M_{22},M_{33}\}$, respectively. Here
$\begin{split}\Lambda_{1}(t)=&(1-\gamma)\sqrt{\alpha(1-\alpha)}-\gamma[(1-\alpha)(1-2q)(1-\gamma)+q(1-\gamma q)],\\ M_{11}&=\frac{(\lambda_{1}-\lambda_{3})^{2}(\lambda_{2}+\lambda_{3})+a^{2}(\lambda_{2}-\lambda_{3})^{2}(\lambda_{1}+\lambda_{3})}{(\lambda_{1}+\lambda_{3})(\lambda_{2}+\lambda_{3})(1+a^{2})}\\ +&\frac{(\lambda_{1}-\lambda_{4})^{2}(\lambda_{2}+\lambda_{4})+b^{2}(\lambda_{2}-\lambda_{4})^{2}(\lambda_{1}+\lambda_{4})}{(\lambda_{1}+\lambda_{4})(\lambda_{2}+\lambda_{4})(1+b^{2})}=M_{22},\\ &M_{33}=\frac{(\lambda_{3}-\lambda_{4})^{2}}{\lambda_{3}+\lambda_{4}}\cdot\frac{(ab-1)^{2}}{(1+a^{2})(1+b^{2})},\end{split}$ (9)
and $\lambda_{i}$ are the eigenvalues of $\rho(t)$, $a=\frac{\rho_{11}(t)-\rho_{44}(t)-\sqrt{(\rho_{11}(t)-\rho_{44}(t))^{2}+4\rho_{14}^{2}(t)}}{2\rho_{14}(t)}$, $b=\frac{\rho_{11}(t)-\rho_{44}(t)+\sqrt{(\rho_{11}(t)-\rho_{44}(t))^{2}+4\rho_{14}^{2}(t)}}{2\rho_{14}(t)}$.
For $q=1$, we have $\Lambda_{1}(t)=2(1-\gamma)[\sqrt{\alpha(1-\alpha)}-\alpha\gamma]>0$ for all $t$, whenever $\alpha<1/2$ (see Fig. 2). On the other hand, the concurrence vanishes ($\Lambda_{1}(t\rightarrow\infty)<0$) for all $\alpha$ when $q\neq 1$ (see Fig. 1). However, the IP exhibits different behavior compared with the concurrence. In Figs. 2 and 2, it is easy to see that the IP decays monotonically and vanishes asymptotically in both situations ($q=1$ and $q=2/3$), and the IP equals zero all the time for $\alpha=0$ and $\alpha=1$ (pure separable states). In addition, an interesting result is that the dynamics of IP exhibit sudden change under the GAD channel while the quantum discord decays smoothly.
3.3 Depolarizing
Figure 3: Dynamics of (a) concurrence and (c) IP as functions of $\alpha$ and $\gamma$, under depolarizing noise channel.
Depolarizing channel describes a process that takes a state into completely mixed state $\mathbb{1}/2$ with probability $\gamma$ and the state being left untouched with probability $1-\gamma$. The Kraus operators of this channel are $E_{0}=\sqrt{1-3\gamma/4}\mathbb{1}$, $E_{1}=\sqrt{\gamma/4}\sigma_{1}$, $E_{2}=\sqrt{\gamma/4}\sigma_{2}$, $E_{3}=\sqrt{\gamma/4}\sigma_{3}$, where $\gamma$ is defined above.[1] Similarly we assume the initial condition as Eq. (7) and then we have the output density matrix elements:
$\begin{split}\rho_{11}(t)&=(1-\alpha)(1-\gamma)+\gamma^{2}/4,\\ \rho_{22}(t)&=\rho_{33}(t)=(2-\gamma)\gamma/4,\\ \rho_{44}(t)&=1-\rho_{11}(t)-2\rho_{22}(t),\\ \rho_{14}(t)&=\rho_{41}(t)=\sqrt{\alpha(1-\alpha)}(1-\gamma)^{2}.\end{split}$ (10)
By Eq. (1), the explicit expression of concurrence for this state is $C(\rho)=2\max\{0,\alpha(1-\gamma)-(2-\gamma)\gamma/4\}$. It is easy to find that $\alpha(1-\gamma)-(2-\gamma)\gamma/4<0$ when $t\rightarrow\infty$ for all $\alpha$ (see Fig. 3). On the other hand, IP does not disappear in a finite time (see Fig. 3). Here the expression of IP can be obtained as the same procedure above.
3.4 dephasing and amplitude damping
Now we consider that two distinct reservoirs (GAD and phase damping) interact with a two-qubit system individually. Given the density matrix of the initial state, the density matrix elements decay at the sum of the amplitude and phase rates, respectively. For simplicity, we assume equal decay rates ($\Gamma$) for both channels, $q=1$ and $\gamma=1-e^{-\Gamma t}$. For the case of the initial condition as Eq. (7), neither amplitude (for $\alpha<1/2$) nor dephasing channel (for all $\alpha$) alone can induce the sudden death of entanglement. However, the dynamics of entanglement are strikingly different when these two kinds of noises are applied together. The concurrence here appears ESD for almost all $\alpha$, as we can see in Fig. 4. This phenomenon of nonadditivity of decoherence while the relaxation rate is additive was first indicated by Yu and Eberly.[23] But, the IP decays asymptotically and smoothly at the same decoherence channel (see Fig. 4), which implies that the additivity of the decoherence channel is still valid for IP.
Figure 4: Dynamics of (a) concurrence and (c) IP as functions of $\alpha$ and $\gamma$, under dephasing and generalized amplitude damping (q=1) channels acted simultaneously.
4 Conclusion
In summary, we investigate the dissipative dynamics of IP for a two-qubit system coupling to independent noisy channels. We observed that by choosing the initial conditions properly, the entanglement of the quantum system suddenly disappears while the IP vanishes only in asymptotic time. This observation indicates that IP is more robust than entanglement and can be considered a better measure of the quantum resources used for some quantum information and computation processes.
Acknowledgments
F. L. Z. was supported by the National Natural Science Foundations of China (Grants Nos. 11675119). J. L. C was supported by the National Natural Science Foundations of China (Grants Nos. 12275136 and 12075001). D. Z. was supported by the Nankai Zhide Foundations.